When you throw a coin n times, Probability Head is 2/3, Tail is 1/3. Find a probability that the pattern HH will never come out.(in nth term)
Answer #1

P(head) = 2/3

P(Tail) = 1/3

P(HH...pattern in nth term) = (2/3) * (2/3) * (2/3) .......n times

= (2/3)^n

P(HH pattern will never come out) = 1 - (2/3)^n